OpenStudy (anonymous):

Ready for a True Challenge? Solve: (9x^2-4)^4-10(9x^2-4)^2+9=0

OpenStudy (anonymous):

\[(9x^2 - 4)^4 - 10(9x^2 - 4) 2 + 9 = 0\] look like that?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

actually the last pary should be a Square^2 +9

OpenStudy (anonymous):

\[(9x^2 - 4)^4 - 10(9x^2 - 4)^2 + 9 = 0\] got it

OpenStudy (anonymous):

Yes that is it :) thank you!! sorry i am new to this and learning how to write the equations

OpenStudy (anonymous):

\[(9x^2 - 4)^4\] is \[(9x^2 - 4)(9x^2 - 4)(9x^2 - 4)(9x^2 - 4)\] right?

OpenStudy (anonymous):

yes...thats as far as i got..

OpenStudy (anonymous):

At first i thought oh this is easy x is 0 because it seems everything cancels itself out...but sill confused

OpenStudy (goformit100):

Well done @Tameiks315

OpenStudy (anonymous):

i got it right? lol

OpenStudy (anonymous):

I wrote the problem incorrectly it should be \[(9x^2-4)^4-10(9x^2-4)^2+9\]

OpenStudy (anonymous):

and yes =0

OpenStudy (anonymous):

@lgbasallote please help :)

OpenStudy (lgbasallote):

the first step is to let a = 9x^2 - 4 so your equation becomes \[a^4 - 10a^2 + 9\] okay?

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

now... i let b = a^2 so this becomes \[\huge b^2 - 10b + 9\] do you follow?

OpenStudy (anonymous):

yes you are factoring it out..i tried but definately missed a step

OpenStudy (lgbasallote):

now.... i break down the middle term \[\huge b^2 - 9b - b + 9\]do you know how to factor this out?

OpenStudy (anonymous):

do you mind breaking it down please...

OpenStudy (anonymous):

i have several problems such as this so if i can have one as a true worked out example it would help me greatly

OpenStudy (lgbasallote):

let me show you an example \[\huge x^2 + 2x + 1\] \[\huge \implies x^2 + x + x + 1\] \[\huge \implies (x^2 + x) + (x+1)\] \[\huge \implies x(x + 1) + (x+1)\] \[\huge \implies (x+1)(x+1)\] got it now?

OpenStudy (anonymous):

i think so...\[(b-3)(b-3)=0\]

OpenStudy (anonymous):

this is where i get stuck because b would equal 3?

OpenStudy (lgbasallote):

it would if the answer was (b-3)(b-3) = 0...but it's not

OpenStudy (lgbasallote):

\[\huge (b^2 - 9b - b + 9)\] \[\huge \implies b(b - 9) - (b-9)\] so what do you think is the factored form?

OpenStudy (anonymous):

i am still stuck :/ b=-9, b=9

OpenStudy (lgbasallote):

what? how did you get that?

OpenStudy (lgbasallote):

what do you get if you factor out b- 9 from what i last wrote?

OpenStudy (anonymous):

i totally bypassed the b in fron of the (b-9)-(b-9) for some reason when my screen updated the equations come over funny for some reason it sayd huge and has brackets

OpenStudy (anonymous):

break out (9x^2 -4)^2 first. (81x^4 -72x^2+16) then i multiplied that by -10 -810x^4+720x^2-160+9 -810x^4+720x^2-151 now, remember that (9x^2 -4)^2 is (81x^4 -72x^2+16) right? so square that to get your equation. (81x^4 -72x^2+16)(81x^4 -72x^2+16) 6561x^8-1164x^6+7776x^4-2304x^2+256 so now you can add that term with your first term just as the equation calls for (9x^2-4)^2+(9x^2-4)^4 6561x^8-1164x^6+7776x^4-2304x^2+256 + -810x^4+720x^2-151 ---------------------------------------- 6561x^8-1164x^6+6696x^4-1584x^2+105 = 0 factor 3(3x-1)(3x+1)(3x^2-1)(9x^2-7)(9x^2-5)=0 so 3x-1=0 3x+1=0 3x^2-1=0 9x^2-7=0 9x^2-5=0 thus your answers are x = -1/3 x=1/3 x=1/sqrt(3) x=-1/sqrt(3) and x=sqrt(5)/3

OpenStudy (anonymous):

How did you get the 72x?